OpenStudy (anonymous):

-7u to 2nd power x4v to 2nd power x 5uv to the 2nd power. how do you simplify?

OpenStudy (anonymous):

are you adding those?

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

(-7u^2x)(4v^2x)(5uv^2)?

OpenStudy (anonymous):

hm

OpenStudy (anonymous):

how about it this way -7u^2 x 4v^2 x 5uv^2 is that easier to read?

OpenStudy (anonymous):

in order to simplify, first factor out the constants so (-7)(4)(5)(u^2)(v^2)(uv^2). then, since (u^2)(u)=u^3 and (v^2)(v^2)=v^4 you get (-120)(u^3)(v^4)

OpenStudy (anonymous):

ok thanks